OpenStudy (anonymous):

Please help & explain how to do it!! Which of the following is an equivalent form of the compound inequality −22 > −5x − 7 −3? −5x − 7 < −22 and −5x − 7 −3 −5x − 7 > −22 and −5x − 7 −3 −5x > −22 and −7 −3 −5x − 7 < −22 and −5x − 7 −3

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

are you sure there is no < or > between the -7 and the -3

OpenStudy (anonymous):

oh! It's the > with the line under it.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[-22>-5x-7 \ge-3\]

OpenStudy (anonymous):

yes. Sorry about the mistake..

OpenStudy (anonymous):

ok it has to be answer A or D; would you mind rewriting the -5x - 7 -3 on the first and last answer, or I can tell

OpenStudy (anonymous):

btw, i think that the answer has to have the word "or" not "and"

OpenStudy (anonymous):

okay, and yes none of the answers have and or or. weird not thank you!

OpenStudy (anonymous):

i get for an answer\[-22>-5x-7 \ or\ -5x-7 \ge-3\]

OpenStudy (anonymous):

which would produce the solution \[(-\infty,-4/5) \cup(3,\infty)\]or graphically